40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 10 00:09:33 2012, kiwamu, Update, LSC, spike hunting in REFL33 
    Reply  Tue Jan 17 01:02:07 2012, kiwamu, Update, LSC, glitch hunting in REFL RFPDs : strange PRMsingle_bounce.png
       Reply  Tue Jan 17 02:27:49 2012, kiwamu, Update, LSC, fringe tests : all the suspensions are innocent 
       Reply  Tue Jan 17 13:47:40 2012, kiwamu, Update, LSC, dirty steering mirror in the REFL path 
          Reply  Wed Jan 18 12:36:26 2012, kiwamu, Update, LSC, wiped a steering mirror on the REFL path DSC_3861_small.jpgDSC_3863_small.jpg
       Reply  Tue Jan 17 19:07:47 2012, rana, Update, LSC, glitch hunting in REFL RFPDs : strange 
       Reply  Wed Jan 18 14:28:36 2012, kiwamu, Update, LSC, estimation of optical length between PRM and scattering object calibration.pngscattering.png
          Reply  Wed Jan 18 16:31:10 2012, kiwamu, Update, LSC, estimation of optical length between PRM and scattering object ND1atten.png
Message ID: 6202     Entry time: Tue Jan 17 01:02:07 2012     In reply to: 6183     Reply to this: 6203   6206   6208   6211
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: glitch hunting in REFL RFPDs : strange 

A very strange thing is going on.

The REFL11 and REFL55 demod signals show high frequency noise depending on how big signals go to the POS actuator of PRM.

This noise shows up even when the beam is single-bounced back from PRM ( the rest of the suspensions are misaligned) and it's very repeatable.

Any idea ?? Am I crazy ?? Is PRM making some fringes with some other optics ??

 

 

(background)

 The most annoying thing in the central part locking is glitches showing up in the LSC error signals (#6183).
The symptom is that when the motion in PRCL at 3 Hz becomes louder, somehow we get glitches in both the MICH and PRCL error signals.
In the frequency domain, those glitches are mostly contribute to a frequency band of about 30 - 100 Hz.
Last Thursday Koji and I locked the half PRM (PRMI with either ITMX or ITMY misaligned) to see if we still have the glitches in this simpler configuration.
Indeed there were the same kind of glitches --a loud 3 Hz motion triggers the glitches.
It was shown particularly in the REFL11 signal but not so much in the REFL33 while AS55 didn't show any glitches.
 
 
(Still glitches even in the single bounce beam)
   We were suspecting some kind of coupling from a beam jitter, so that the 3 Hz motion somehow brings the beam spot to a bad place somewhere in the REFL paths.
I misaligned all the suspensions except for PRM such that the beam directly bounces back from PRM and go to the REFL port.
Indeed there still were glitches in the REFL11 and REFL55 demod signals. It showed up once per 30 sec or so and pushes up the noise floor around 30 - 100 Hz.
There might be a little bit of glitches also in the REFL33, but the ADC noise floor and the expected glitch noise level were comparable and hence it was difficult to see the glitches in REFL33.
 

(Glitch is related to the PRM POS actuation)

 In the single-bounce configuration I started shaking the PIT and YAW motions of PRM at 3 Hz using the realtime LOCKIN oscillator to see if I can reproduce the glitches.
However no significant glitches were found in this test.
Then I started shaking the POS instead of the angular DOFs, and found that it causes the glitches.
At this point it didn't look like a glitch any more, it became more like a stationary noise.
 
 The attached screen shot is the noise spectrum of the REFL11_I.
The red curve is the one taken when I injected the 3 Hz excitation in POS by the LOCKIN oscillator.
The excitation is at 3 Hz with an amplitude of 1000 counts.
As a comparison I plotted the same spectrum when no excitation was injected and it is plotted in pink.
 

 PRMsingle_bounce.png

It seems there is a cut off frequency at 100 Hz.
This frequency depends on the amplitude of the excitation -- increasing the amplitude brings the cut off frequency higher.
This noise spectrum didn't change with and without the oplevs and local damping.
 

(Possible scenario)

A possible reason that I can think of right now is : PRM is interfering with some other optics for some reason.
But if it's true, why I didn't see any fringes in the AS demod signals in the half PRM configuration ?
 

Quote from #6183

 We tried to figure out what is causing spikes in the REFL33 signal, which is used to lock PRCL.

No useful information was obtained tonight and it is still under investigation.

ELOG V3.1.3-